# Revisiting Semi-Supervised Learning in the Era of Foundation Models

### v3

@@ Abstract
footnotetext: *Equal contribution.footnotetext: Our code is available at https://github.com/OSU-MLB/SSL-Foundation-Models.
@@ 1 Introduction
Key insights. Our empirical results highlight two main findings. First, fine-tuning VFMs with representative SSL algorithms offers limited advantages over using only labeled data for fine-tuning. Second, parameter-efficient fine-tuning (PEFT) mai2024lessonslearnedunifyingempirical ; vpetlbench2024 —which updates a small subset of parameters or adds lightweight learnable modules while keeping the VFM largely frozen—consistently yields substantial performance gains, regardless of the learning paradigm.
@@ 2 Related Works
Vision Foundation Models. Vision Transformers (ViT) dosovitskiy2021imageworth16x16words pre-trained on massive amounts of data, has become indispensable to modern AI development awais2025foundation ; mai2025ava ; gu2025bioclip . These models, often referred to as vision foundation models (VFMs), have demonstrated superior performance on a wide range of tasks. For example, CLIP-ViT radford2021learningtransferablevisualmodels trained with millions of image-text pairs shows an unprecedented zero-shot capability, robustness to distribution shifts, and serves as the encoders for various powerful generative models rombach2022high ; liu2024visual . Meanwhile, DINOv2 oquab2024dinov2learningrobustvisual pre-trained with self-supervised objectives on extensive sets of well-curated images effectively captures fine-grained localization features. Moreover, it is increasingly becoming a common sense that strategically fusing multiple VFMs can unlock synergistic gains mai2024lessonslearnedunifyingempirical ; tong2024eyes , leading to state-of-the-art performance across tasks from visual question answering to object detection.
@@
Table 1: Comparison of supervised linear probing performance between popular SSL datasets (CIFAR, Food101) and our benchmark, where N denotes the number of labeled samples per class. While frozen VFMs already excel in standard SSL datasets, they struggle in our benchmark—comprising diverse tasks, domains, and sizes—underscoring the potential of SSL for unleashing the full potential of VFMs.
@@ 3.2 A Comprehensive SSL Image Classification Benchmark
Dataset & Regime We choose six classification datasets from VTAB zhai2020largescalestudyrepresentationlearning , spanning remote sensing, medical imaging, synthetic data, and autonomous driving tasks. To test the robustness of SSL methods, we vary the number of labels per class and employ linear probing as our performance metric, selecting shot counts that keep each task sufficiently challenging for frozen VFMs. These settings impose substantial difficulty on frozen VFMs (Table 1), underscoring the importance of SSL for unlocking their full potential. A summary of the datasets appears in Table 2, with full details provided in Appendix A. Together, this benchmark offers a more diverse and comprehensive foundation‐model‐era evaluation of semi‐supervised image classification.
@@ 3.3 Fair Hyperparameter Tuning for SSL
However, ericsson2023betterpracticesdomainadaptation also revealed that no single criterion could reliably select suitable hyperparameters across all scenarios. Motivated by this finding, we propose integrating seven unsupervised criteria—five derived from features, AMI vinh2009information , ARI halkidi2002cluster , V-Measure rosenberg2007v , FMI fowlkes1983method and BNM cui2020towards , and two from logits, RankMe garrido2023rankmeassessingdownstreamperformance and CHI calinski1974dendrite —for more robust tuning. Concretely, for each hyperparameter configuration and its corresponding model, we compute all seven criteria using a held-out unlabeled validation set $\mathcal{V}\coloneqq\{{\bm{x}}^{v}_{i}\}_{i=1}^{|\mathcal{V}|}$. Next, we rank every hyperparameter configuration based on each criterion and choose the one achieving the lowest average rank across all. We provide formal definitions and detailed procedures in section˜C.1 and the effectiveness of the proposed method in section˜B.7. By eliminating reliance on held-out labeled validation sets, our method mitigates data leakage and promotes a more practical and fair tuning protocol for SSL.
@@ 5.1 Self-training
In our context, a VFM fine-tuned with PEFT using labeled $\mathcal{L}$ serves as the current model to generate pseudo-labels for $\mathcal{U}$, which are then used to further fine-tune the VFM.
@@ 6.1 Performance Comparison
We summarize the performance comparison between ST, PET, V-PET and existing SSL methods across all the 12 settings in Table 3. Overall, V-PET achieve superior performance compared with others, as shown in the Average column. To further analyze the results, Figure 6 visualize the ranking frequency, where each matrix entry $(i,j)$ indicates how often method $i$ is ranked $j$-th across 12 settings. We then compute the mean rank (number in the bracket) of each method and sort them accordingly. Although V-PET does not claim first place in every single setting, it secures the top rank most frequently, showcasing the effectiveness of pseudo-label ensemble and establishing it as a simple yet powerful baseline in the VFM era.
@@
LoRA Labeled Only 57.9% 64.1% 60.7% 66.8% 59.5% 72.1% 35.4% 42.2% 35.2% 50.8% 60.1% 64.3% 55.7%
Fixmatch 56.8% 68.6% 62.2% 72.8% 47.6% 77.4% 32.5% 39.1% 32.2% 42.6% 59.8% 52.6% 53.7%
FlexMatch 56.5% 67.0% 66.4% 71.9% 51.2% 78.4% 32.8% 38.7% 36.9% 75.4% 44.6% 55.0% 56.2%
SoftMatch 59.8% 68.0% 66.3% 70.9% 78.7% 83.7% 42.1% 42.9% 29.8% 71.2% 42.9% 60.5% 59.7%
FineSSL 63.1% 69.1% 56.3% 61.8% 74.2% 83.5% 35.0% 30.4% 33.6% 38.1% 53.7% 48.2% 53.9%
ST 58.9% 63.8% 49.2% 52.7% 64.1% 77.4% 35.7% 42.8% 36.5% 51.6% 59.8% 64.3% 54.7%
PET 63.4% 68.2% 65.6% 70.8% 67.5% 78.6% 36.4% 42.4% 38.3% 54.6% 60.3% 65.7% 59.3%
V-PET 65.6% 71.7% 67.2% 72.8% 66.6% 77.2% 38.9% 51.1% 36.7% 58.5% 58.0% 62.3% 60.5%
AdaptFormer Labeled Only 61.5% 65.0% 61.1% 68.2% 61.0% 72.8% 34.4% 39.0% 37.8% 48.6% 57.8% 59.1% 55.6%
Fixmatch 61.2% 67.3% 64.6% 71.2% 63.1% 80.5% 29.0% 42.4% 33.1% 44.8% 30.0% 57.2% 53.7%
FlexMatch 60.3% 68.8% 64.3% 70.7% 58.1% 78.1% 31.0% 35.9% 38.5% 52.2% 34.2% 54.2% 53.9%
SoftMatch 61.3% 69.3% 65.3% 70.2% 65.8% 82.7% 38.2% 42.7% 34.5% 56.1% 33.9% 55.1% 56.3%
FineSSL 61.9% 68.1% 58.4% 64.9% 73.9% 84.2% 25.3% 29.8% 26.0% 36.9% 42.2% 47.1% 51.6%
ST 62.0% 67.2% 58.1% 61.3% 68.8% 79.3% 34.4% 39.6% 39.0% 50.3% 58.2% 61.9% 56.7%
PET 63.8% 68.9% 66.8% 71.7% 69.5% 80.2% 37.1% 41.5% 38.4% 55.8% 60.5% 61.7% 59.7%
V-PET 65.7% 71.8% 67.8% 73.2% 67.9% 78.4% 38.9% 51.2% 37.0% 59.0% 58.2% 63.3% 61.0%
@@
LoRA Labeled Only 61.0% 68.4% 59.7% 66.8% 48.6% 63.1% 40.7% 48.0% 35.9% 61.7% 46.7% 71.7% 56.0%
Fixmatch 57.4% 68.4% 56.6% 71.4% 36.6% 66.7% 38.1% 41.2% 35.7% 41.0% 14.8% 42.9% 47.6%
FlexMatch 52.8% 65.4% 55.7% 67.7% 33.3% 72.6% 36.6% 35.8% 37.1% 66.9% 26.2% 33.6% 48.6%
SoftMatch 53.6% 64.0% 60.4% 67.3% 46.6% 79.5% 32.1% 44.5% 33.6% 65.1% 27.4% 27.9% 50.2%
FineSSL 66.9% 72.7% 54.6% 62.1% 69.8% 87.4% 31.0% 24.6% 24.8% 31.7% 31.9% 39.8% 49.8%
ST 52.3% 56.9% 45.4% 50.5% 54.7% 72.4% 42.6% 49.6% 36.3% 62.7% 33.2% 77.8% 52.9%
PET 66.7% 72.2% 58.6% 65.3% 55.8% 70.6% 41.0% 54.3% 33.1% 56.9% 42.6% 63.3% 56.7%
V-PET 67.8% 74.1% 67.1% 73.2% 66.0% 77.0% 39.2% 52.5% 36.8% 59.3% 58.2% 63.7% 61.2%
AdaptFormer Labeled Only 63.0% 67.6% 58.9% 67.0% 47.0% 57.4% 34.4% 53.1% 27.3% 41.8% 52.3% 49.8% 51.6%
Fixmatch 59.5% 70.4% 57.9% 69.7% 26.4% 62.3% 38.7% 45.6% 32.7% 38.9% 19.8% 50.4% 47.7%
FlexMatch 57.1% 65.9% 56.0% 67.5% 51.3% 64.0% 33.4% 40.2% 33.4% 42.2% 34.3% 27.9% 47.8%
SoftMatch 58.7% 66.2% 60.0% 66.7% 53.1% 78.1% 33.5% 32.2% 32.9% 45.0% 24.5% 30.9% 48.5%
FineSSL 64.0% 71.2% 56.0% 63.3% 62.3% 69.7% 17.7% 24.3% 25.8% 30.0% 31.8% 45.0% 46.8%
ST 62.3% 67.0% 59.0% 67.1% 54.1% 66.2% 34.9% 56.0% 27.4% 43.1% 43.9% 48.7% 52.5%
PET 67.2% 73.0% 63.5% 70.9% 56.2% 72.0% 40.2% 55.5% 33.1% 58.0% 45.2% 60.9% 58.0%
V-PET 68.2% 74.2% 67.7% 73.4% 66.3% 78.0% 39.4% 52.5% 36.8% 60.0% 59.2% 61.7% 61.5%
Table 3: Performance comparison of baselines, existing SSL approaches, and our proposed methods on six diverse datasets (12 settings). The best result within each PEFT and VFM is highlighted in bold. We report the full results in section˜B.2.
@@ 6.2 Scalability of Ensembling
Figure 5: Scaling analysis for ensembling more PEFT methods and VFMs. Starting from LoRA fine-tuning, performance improves with diverse pseudo-labels, but gain diminishes with more ensemble sources.
@@
Figure 6: Ranking frequency across SSL methods by the proposed benchmark. The number in $(i,j)$ indicates the frequency of method $i$ is ranked $j$-th across 12 settings. The number in bracket indicates the average rank.
@@ 7 Conclusion
As the machine learning community increasingly standardizes pre-trained foundation models as backbones across various applications, it becomes crucial to assess whether machine learning algorithms developed prior to the foundation model era remain effective. Specifically, given that vision foundation models (VFM) are pre-trained on millions of images, it remains unclear whether they would still benefit from abundant unlabeled data in conjunction with limited labeled data in a semi-supervised learning (SSL) setting. To this end, we conduct a comprehensive and systematic empirical study, revealing that VFMs can still benefit from SSL but require a specifically tailored approach. In particular, we propose a simple yet highly effective self-training algorithm that ensembles pseudo-labels from multiple VFMs and parameter-efficient fine-tuning (PEFT) techniques. Overall, our study offers actionable insights into SSL with VFMs and lays the groundwork for more scalable and practical semi-supervised learning in the era of foundation models.
@@ B.6 The Effectiveness of VFMs Ensemble
B.7 The Effectiveness of Hyperparameter Tuning
Table 11: Hyperparameter tuning comparison: our method versus baselines (random or chosen by one criterion). The absolute errors from oracle test accuracy (smaller the better) are reported averaged across 12 datasets, 5 SSL baselines, and 2 VFMs trained on 2 PEFT. The reported values are in the format: mean $\pm$ standard deviation.
RankMe AMI ARI V-Measure FMI CHI BNM Random Ours
[…]
We begin by investigating whether the proposed hyperparameter tuning strategy accurately gauges SSL methods’ performance. To illustrate its advantages, we compare it with two baselines, both measured via the absolute difference from the oracle test performance. Specifically, the oracle test performance is defined as the highest accuracy across the entire hyperparameter search space; each method selects a hyperparameter configuration, and we record how far its resulting accuracy is from this optimum. By taking the absolute value of this difference, we quantify how closely each tuning method approximates the best possible result.
@@ C.2 Validators
Adjusted Rand Index (ARI) [30] ARI is a widely used metric in cluster analysis by measuring the similarity between two clustering solutions. Given $RI$ is the Rank Index between the true and predicted clustering, ARI is defined by:

### v4

@@ Abstract
footnotetext: Our code is available at https://github.com/OSU-MLB/SSL-Foundation-Models.
@@ 1 Introduction
Key insights. Our empirical results highlight two main findings. First, fine-tuning VFMs with representative SSL algorithms offers limited advantages over using only labeled data for fine-tuning. Second, parameter-efficient fine-tuning (PEFT) mai2024lessonslearnedunifyingempirical ; vpetlbench2024 ; sun2025strongermixturelowrankexperts —which updates a small subset of parameters or adds lightweight learnable modules while keeping the VFM largely frozen—consistently yields substantial performance gains, regardless of the learning paradigm.
@@ 2 Related Works
Vision Foundation Models. Vision Transformers (ViT) dosovitskiy2021imageworth16x16words pre-trained on massive amounts of data, have become indispensable to modern AI development awais2025foundation ; mai2025ava ; gu2025bioclip . These models, often referred to as vision foundation models (VFMs), have demonstrated superior performance on a wide range of tasks. For example, CLIP-ViT radford2021learningtransferablevisualmodels trained with millions of image-text pairs shows an unprecedented zero-shot capability, robustness to distribution shifts, and serves as the encoders for various powerful generative models rombach2022high ; liu2024visual . Meanwhile, DINOv2 oquab2024dinov2learningrobustvisual pre-trained with self-supervised objectives on extensive sets of well-curated images effectively captures fine-grained localization features. Moreover, it is increasingly becoming common sense that strategically fusing multiple VFMs can unlock synergistic gains mai2024lessonslearnedunifyingempirical ; tong2024eyes , leading to state-of-the-art performance across tasks from visual question answering to object detection.
@@
Table 1: Comparison of supervised linear probing performance (%) between popular SSL datasets (CIFAR, Food101) and our benchmark, where N denotes the number of labeled samples per class. While frozen VFMs already excel in standard SSL datasets, they struggle in our benchmark—comprising diverse tasks, domains, and sizes—underscoring the potential of SSL for unleashing the full potential of VFMs.
@@ 3.2 A Comprehensive SSL Image Classification Benchmark
Dataset & Regime Following the VTAB protocol zhai2020largescalestudyrepresentationlearning , we select six classification datasets covering the three VTAB categories— Natural , Specialized , and Structured . Specifically, we choose two datasets from each category: DTD and SUN397 from Natural , RESISC45 and Retinopathy from Specialized , and CLEVR-C and KITTI from Structured . These datasets span diverse domains, including texture recognition, scene understanding, remote sensing, medical imaging, synthetic reasoning, and autonomous driving. To evaluate the robustness of SSL methods, we vary the number of labeled samples per class and adopt linear probing as the evaluation protocol, with shot counts chosen to keep each task sufficiently challenging for frozen VFMs. This configuration imposes substantial difficulty on frozen representations (Table 1), underscoring the necessity of SSL for unlocking their full potential. A summary of the datasets appears in Table 2, and detailed descriptions are provided in Appendix A. Together, this benchmark offers a more diverse and comprehensive foundation‐model‐era evaluation of semi‐supervised image classification.
@@ 3.3 Fair Hyperparameter Tuning for SSL
However, ericsson2023betterpracticesdomainadaptation also revealed that no single criterion could reliably select suitable hyperparameters across all scenarios. Motivated by this finding, we propose integrating seven unsupervised criteria—five derived from features, AMI vinh2009information , ARI halkidi2002cluster , V-Measure rosenberg2007v , FMI fowlkes1983method and BNM cui2020towards , and two from logits, RankMe garrido2023rankmeassessingdownstreamperformance and CHI calinski1974dendrite —for more robust tuning. Concretely, for each hyperparameter configuration and its corresponding model, we compute all seven criteria using a held-out unlabeled validation set $\mathcal{V}\coloneqq\{{\bm{x}}^{v}_{i}\}_{i=1}^{|\mathcal{V}|}$. Next, we rank every hyperparameter configuration based on each criterion and choose the one achieving the lowest average rank across all. We provide formal definitions and detailed procedures in section˜C.1 and the effectiveness of the proposed method in section˜6.3. By eliminating reliance on held-out labeled validation sets, our method mitigates data leakage and promotes a more practical and fair tuning protocol for SSL.
@@ 5.1 Self-training
In our context, a VFM fine-tuned with PEFT using labeled dataset $\mathcal{L}$ serves as the current model to generate pseudo-labels for $\mathcal{U}$, which are then used to further fine-tune the VFM.
@@ 5.2 Ensembling of multiple PEFT and VFMs
Time Efficiency of Ensembling. Among steps (a)(d) in algorithm 1, step (d) dominates the wall-clock time since the pseudo-labeled set is orders of magnitude larger than the labeled one. Although our method ensembles $N\times M$ models, the overall time overhead remains marginal: step (a) is fast owing to the small labeled set, and steps (b) and (c) are computationally light. For instance, for V-PET, the estimated runtime is only about 1.16× that of other SSL baselines.
@@ 6.1 Performance Comparison
We summarize the performance comparison between ST, PET, V-PET and existing SSL methods across all the 12 settings in Table 3. Overall, V-PET achieves superior performance compared with others, as shown in the Average column. To further analyze the results, Figure 5 visualize the ranking frequency, where each matrix entry $(i,j)$ indicates how often method $i$ is ranked $j$-th across 12 settings. We then compute the mean rank (number in the bracket) of each method and sort them accordingly. Although V-PET does not claim first place in every single setting, it secures the top rank most frequently, showcasing the effectiveness of pseudo-label ensemble and establishing it as a simple yet powerful baseline in the VFM era.
@@
LoRA Labeled Only 57.9 64.1 60.7 66.8 59.5 72.1 35.4 42.2 35.2 50.8 60.1 64.3 55.7
Fixmatch 56.8 68.6 62.2 72.8 47.6 77.4 32.5 39.1 32.2 42.6 59.8 52.6 53.7
FlexMatch 56.5 67.0 66.4 71.9 51.2 78.4 32.8 38.7 36.9 75.4 44.6 55.0 56.2
SoftMatch 59.8 68.0 66.3 70.9 78.7 83.7 42.1 42.9 29.8 71.2 42.9 60.5 59.7
FineSSL 63.1 69.1 56.3 61.8 74.2 83.5 35.0 30.4 33.6 38.1 53.7 48.2 53.9
ST 58.9 63.8 49.2 52.7 64.1 77.4 35.7 42.8 36.5 51.6 59.8 64.3 54.7
PET 63.4 68.2 65.6 70.8 67.5 78.6 36.4 42.4 38.3 54.6 60.3 65.7 59.3
V-PET 65.6 71.7 67.2 72.8 66.6 77.2 38.9 51.1 36.7 58.5 58.0 62.3 60.5
AdaptFormer Labeled Only 61.5 65.0 61.1 68.2 61.0 72.8 34.4 39.0 37.8 48.6 57.8 59.1 55.6
Fixmatch 61.2 67.3 64.6 71.2 63.1 80.5 29.0 42.4 33.1 44.8 30.0 57.2 53.7
FlexMatch 60.3 68.8 64.3 70.7 58.1 78.1 31.0 35.9 38.5 52.2 34.2 54.2 53.9
SoftMatch 61.3 69.3 65.3 70.2 65.8 82.7 38.2 42.7 34.5 56.1 33.9 55.1 56.3
FineSSL 61.9 68.1 58.4 64.9 73.9 84.2 25.3 29.8 26.0 36.9 42.2 47.1 51.6
ST 62.0 67.2 58.1 61.3 68.8 79.3 34.4 39.6 39.0 50.3 58.2 61.9 56.7
PET 63.8 68.9 66.8 71.7 69.5 80.2 37.1 41.5 38.4 55.8 60.5 61.7 59.7
V-PET 65.7 71.8 67.8 73.2 67.9 78.4 38.9 51.2 37.0 59.0 58.2 63.3 61.0
@@
LoRA Labeled Only 61.0 68.4 59.7 66.8 48.6 63.1 40.7 48.0 35.9 61.7 46.7 71.7 56.0
Fixmatch 57.4 68.4 56.6 71.4 36.6 66.7 38.1 41.2 35.7 41.0 14.8 42.9 47.6
FlexMatch 52.8 65.4 55.7 67.7 33.3 72.6 36.6 35.8 37.1 66.9 26.2 33.6 48.6
SoftMatch 53.6 64.0 60.4 67.3 46.6 79.5 32.1 44.5 33.6 65.1 27.4 27.9 50.2
FineSSL 66.9 72.7 54.6 62.1 69.8 87.4 31.0 24.6 24.8 31.7 31.9 39.8 49.8
ST 52.3 56.9 45.4 50.5 54.7 72.4 42.6 49.6 36.3 62.7 33.2 77.8 52.9
PET 66.7 72.2 58.6 65.3 55.8 70.6 41.0 54.3 33.1 56.9 42.6 63.3 56.7
V-PET 67.8 74.1 67.1 73.2 66.0 77.0 39.2 52.5 36.8 59.3 58.2 63.7 61.2
AdaptFormer Labeled Only 63.0 67.6 58.9 67.0 47.0 57.4 34.4 53.1 27.3 41.8 52.3 49.8 51.6
Fixmatch 59.5 70.4 57.9 69.7 26.4 62.3 38.7 45.6 32.7 38.9 19.8 50.4 47.7
FlexMatch 57.1 65.9 56.0 67.5 51.3 64.0 33.4 40.2 33.4 42.2 34.3 27.9 47.8
SoftMatch 58.7 66.2 60.0 66.7 53.1 78.1 33.5 32.2 32.9 45.0 24.5 30.9 48.5
FineSSL 64.0 71.2 56.0 63.3 62.3 69.7 17.7 24.3 25.8 30.0 31.8 45.0 46.8
ST 62.3 67.0 59.0 67.1 54.1 66.2 34.9 56.0 27.4 43.1 43.9 48.7 52.5
PET 67.2 73.0 63.5 70.9 56.2 72.0 40.2 55.5 33.1 58.0 45.2 60.9 58.0
V-PET 68.2 74.2 67.7 73.4 66.3 78.0 39.4 52.5 36.8 60.0 59.2 61.7 61.5
Table 3: Performance (%) comparison of baselines, existing SSL approaches, and our proposed methods on six diverse datasets (12 settings). The best result within each PEFT and VFM is highlighted in bold. We report the full results in section˜B.2.
@@ 6.2 Scalability of Ensembling
Figure 5: Ranking frequency across SSL methods by the proposed benchmark. The number in $(i,j)$ indicates the frequency of method $i$ is ranked $j$-th across 12 settings. The number in brackets indicates the average rank, where the higher rank is better.
6.3 The Effectiveness of Hyperparameter Tuning
Table 4: Hyperparameter tuning comparison: our method versus baselines (random or chosen by one criterion). The absolute errors (%) from oracle test accuracy (smaller the better) are reported averaged across 12 datasets, 5 SSL baselines, and 2 VFMs trained on 2 PEFT. The reported values are in the format: mean $\pm$ standard deviation.
RankMe AMI ARI V-Measure FMI CHI BNM Random Ours
[…]
We also investigated whether the proposed hyperparameter tuning strategy accurately gauges SSL methods’ performance. To illustrate its advantages, we compare it with two baselines, both measured via the absolute difference from the oracle test performance. Specifically, the oracle test performance is defined as the highest accuracy across the entire hyperparameter search space; each method selects a hyperparameter configuration, and we record how far its resulting accuracy is from this optimum. By taking the absolute value of this difference, we quantify how closely each tuning method approximates the best possible result.
Figure 6: Scaling analysis for ensembling more PEFT methods and VFMs. Starting from LoRA fine-tuning, performance improves with diverse pseudo-labels, but the gain diminishes with more ensemble sources. Mean Label consistently outperforms Mean Logits and Mean Probabilities.
@@ 7 Conclusion
We conduct a comprehensive and systematic empirical study for SSL in the era of VFMs, revealing that VFMs can still benefit from SSL but require a specifically tailored approach. In particular, we propose a simple yet highly effective self-training algorithm that ensembles pseudo-labels from multiple VFMs and parameter-efficient fine-tuning (PEFT) techniques. Overall, our study offers actionable insights into SSL with VFMs and lays the groundwork for more scalable and practical semi-supervised learning in the era of foundation models.
@@ C.2 Validators
Adjusted Rand Index (ARI) [30] ARI is a widely used metric in cluster analysis by measuring the similarity between two clustering solutions. Given $RI$ is the Rand Index between the true and predicted clustering, ARI is defined by: